# Attention-Based Multi-Scale Temporal Fusion Network for Uncertain-Mode Fault Diagnosis in Multimode Processes

### v1

@@ 4.4.2 Experimental results and analysis
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
@@
The feature visualization results obtained using t-SNE are shown in Fig. 11. Although CNN-LSTM achieves a Micro F1 score of 0.9997, a comparison of Fig. 11(b) and (e) clearly shows that the features extracted by CNN-LSTM exhibit weaker intra-class compactness and inter-class separability compared to those extracted by AMTFN. From Fig. 11(a), (c) and (d), it can be seen that the features extracted by these models form poorly separable clusters. In contrast, AMTFN achieves promising results, with features of samples from the same class exhibiting regional distributions and clear inter-class distances.
Fig. 11.: Feature map visualization results via t-SNE.
The parameter number, training time per epoch, and testing time per epoch for each model are presented in Table 14. The experimental results are similar to those obtained in TE process. This demonstrates that AMTFN is sufficiently lightweight, proving its advantage in practical applications.
Table 14: Parameter number, training time and test time of different methods on the three-phase flow facility dataset.
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFN
@@ 5 Conclusion
This article proposed a model named AMTFN for uncertain-mode fault diagnosis. The AMTFN stands out by focusing on critical time points with higher cross-mode shared information, significantly enhancing its ability to extract domain-invariant features. This improvement leads to superior performance in uncertain-mode fault diagnosis tasks. The proposed method was evaluated on the TE process and the three-phase flow facility. The experiments demonstrate that AMTFN exhibits significant superiority in terms of fault diagnosis performance, visualization, and lightweight design.
However, uncertain-mode fault diagnosis has certain limitations. It assumes that there is no distribution shift between the training and testing datasets. The industrial systems may experience new faults or operate in new modes. Therefore, exploring open-set uncertain-mode fault diagnosis and addressing the generalization problem in uncertain-mode fault diagnosis could be necessary future directions.

### v2

@@ 4.4.2 Experiment results
DCNN CNN-LSTM IPO-ViT MGAMN AMTFNet
@@
The feature visualization results obtained using t-SNE are shown in Fig. 11. Although the Micro F1 score of CNN-LSTM reaches 0.9997, the comparison of Fig. 11(b) and (e) clearly shows that the features extracted by CNN-LSTM exhibit weaker cohesion of same-class and divergence among different classes compared to those extracted by AMTFNet. The results shown in Fig. 11(a), (c) and (d) demonstrate that the features extracted by these models form poorly separable clusters. In contrast, AMTFNet achieves promising results, with features of samples from the same class exhibiting regional distributions and clear inter-class distances.
Fig. 11.: T-SNE visualization results on the TPFF dataset.
The parameter number, training time per epoch, and testing time per epoch for each model are presented in Table 14. The results are consistent with those obtained in TE process. This demonstrates that AMTFNet is sufficiently lightweight, proving its advantage in practical applications.
Table 14: Experiment results for scalability comparison on the TPFF dataset.
@@
DCNN CNN-LSTM IPO-ViT MGAMN AMTFNet
@@ 5 Conclusion
This article proposed a model named AMTFNet for uncertain-mode fault diagnosis. The AMTFNet stands out by focusing on key temporal moments that exhibit richer cross-mode information, significantly enhancing its ability to extract domain-invariant features. This improvement leads to superior performance in uncertain-mode fault diagnosis tasks. The experiments on two datasets demonstrate that AMTFNet exhibits significant superiority in terms of fault diagnosis performance, visualization, and lightweight design.
However, uncertain-mode fault diagnosis has certain limitations. It assumes that there is no discrepancy in data distribution between the training and test datasets. The industrial systems may experience new faults or operate in new modes. Therefore, exploring open-set uncertain-mode fault diagnosis and addressing the generalization problem in uncertain-mode fault diagnosis could be necessary future directions.